The length of a rectangle varies inversely as the width. When the length is 20 inches, the width is 2 inches. When the length is 5 inches, what is the width
Answer by math_helper(2461)


Since the length and width vary inversely: L*W = k
When length is 20, width is 2 ==> 20*2 = 40 (= k)
So when the length is 5 the width will be 5*W = 40 —> W = 40/5 = 8

Ans: The width is 8in when the length is 5in.